OpenStudy (anonymous):

A bottle and a cap together cost $1.10. The bottle costs $1 more than the cap. How much does each cost?

OpenStudy (anonymous):

Let b = bottle & c = cap We would have to use a system of equations. So, since the bottle and cap together cost $1.10, the first equation is b + c = $1.10 And the bottle IS $1 more than the cap (remember, is always means =) And more means addition. Your second equation is, b = c + 1 Now, you're system of equations is { b + c = $1.10 { b = c + 1 You can use substitution and put in the given value for b. c + 1 + c = $1.10 Then solve for c. x

OpenStudy (anonymous):

Tell me if you got it. x